OpenStudy (anonymous):

the sum of two numbers is 8. If one number is subtracted from the other, the result is -2. Find the number

OpenStudy (anonymous):

3 and 5..... 3+5=8 .... 3-5=-2

OpenStudy (anonymous):

a) x + y = 8 b) x - y = -2 Now, isolate x in equation (a). Then substitute this result into the the value of x in (b). You should now how (b) in terms of only "y". Solve for y. Now you know y. sub it back into (a) or (b) to find the value of "x"

OpenStudy (anonymous):

Let x be the 1st number let y be the 2nd number write the 2 equations x+y=8 x-y=-2 then add so it will become 2x=6 therefore x=3 If x is equal to 3 find y so used the 1st equation x+y=8 (3)+y=8 transfer 3 to another side y=8-3 therefore y=5 the two numbers are 3 and 5